OpenStudy (anonymous):

–3x + y = 2 –9x + 3y = 6 What is the solution to the system? 1.No Solutions 2. Infinite Solutions 3.(3, 2) 4.(9, 2)

jimthompson5910 (jim_thompson5910):

–3x + y = 2 –9x + 3y = 6 -3(–3x + y) = -3(2) –9x + 3y = 6 9x - 3y = -6 –9x + 3y = 6 Add the equations 9x - 3y = -6 –9x + 3y = 6 -------------- 0x + 0y = 0 which leads to 0 = 0 which is always true. So there are an infinite number of solutions.